# Age-Optimal Sampling and Routing under Intermittent Links and Energy Constraints

### v1

@@ VIII Simulation Results
This section presents simulation results for practical scenarios to validate the analytical findings and evaluate the performance of our proposed algorithm.
VIII-A Comparing Benchmarks
@@
In this subsection, we refer to our designed jointly optimal sampling and routing policy as the “optimal policy” and evaluate its performance against the following benchmark policies:
@@
$\bullet$ Minimum Average Delay Routing with AoI-Optimal Sampling (MAD-Optimal): This policy always selects the route with the minimum average delay over the set of available routes at each instance. Given this selection, a modified version of the ReAVI with the minimization over the routing options is implemented to find the AoI-optimal waiting strategy and minimize the long-term average AoI.
@@
$\bullet$ Minimum Average Delay Routing with Zero-Wait Sampling (MAD-Zero Wait): This policy always selects the route with the minimum average delay over the set of available routes at each instance. It is combined with a zero-wait strategy, where a new packet is sampled and transmitted immediately upon the delivery of the previous packet [^3] Zero-wait policy [sun2017update] is work-conserving, hence, it achieves maximum throughput on any given route.. The long term average AoI achieved by this policy can be analytically calculated.
@@
$\quad\quad\ +\left(\frac{\mu_{2}}{2}+\frac{\sigma_{2}^{2}}{2\mu_{2}}\right)\left(p-p^{2}\right)\mu_{2}$
$\quad\quad\ +\left(\frac{\mu_{3}}{2}+\frac{\sigma_{3}^{2}}{2\mu_{3}}\right)p\mu_{3}.$
@@
$\bullet$ Minimum Delay Variance Routing with Zero-Wait Sampling (MDV-Zero Wait): This policy consistently selects the route with the lowest delay variance. It is combined with a zero-wait strategy.
$\bullet$ Minimum Delay Variance Routing with AoI-Optimal Sampling (MDV-Optimal): This policy always selects the route with the minimum delay variance and follows the AoI-optimal waiting strategy as outlined in [sun2017update, Theorem 4].
VIII-B Satellite-Terrestrial Integrated Routes
@@
VIII-B1 LEO Satellite Routes with Stochastic Delays
For $l\in\mathcal{N}_{\text{Sat}}$, the delay is modeled by a log-normal distribution, characterized by the following probability density function [peebles2001probability]:
@@ VIII-B1 LEO Satellite Routes with Stochastic Delays
VIII-B2 Terrestrial Routes with Stochastic Delays
If $l\in\mathcal{N}_{\text{Ter}}$, we leverage the gamma distribution to simulate the statistics of delay $y$, where the probability density function is given by [peebles2001probability]:
@@ VIII-B2 Terrestrial Routes with Stochastic Delays
VIII-C Parameter Settings
Figure 4: AoI vs. $E_{\max}$
Figure 7: Simulation results of systems with $N=2$ and $N=3$.
TABLE I: Simulation Parameters $\boldsymbol{p}=\boldsymbol{1},E_{\max}=\infty$
Route Route 11 Route 22 Route 33
Distribution Log-normal Gamma Gamma
Parameters $(\mu_{1},\sigma_{1})$ $(\mu_{2},\sigma_{2})$ $(\mu_{3},\sigma_{3})$
Fig. 7 (a) $(3.4,[0,3])$ $(0.7,5)$
Fig. 7 (b) $([0.7,5.2],2)$ $(0.7,5)$
Fig. 7 (c) $(2.4,[0,3])$ $(1.2,3)$ $(0.7,3.4)$
Fig. 7 (d) $(2.4,0.7)$ $(1.2,3)$ $([0.7,1.2],3.4)$
Fig. 8 $(2.4,0.7)$ $(1.2,3)$ $(0.7,3.4)$
TABLE II: Simulation Parameters $E_{\max}=\infty$
@@ VIII-C Parameter Settings
We first consider a scenario where there are $2$ available routes that are always available with the varied energy constraint $E_{\max}\in[1,5]$. The routes has the delay statistics: Log-normal distribution, $\mu_{1}=5,\sigma_{1}=1$; Gamma distribution, $\mu_{2}=1,\sigma_{2}=7.3$. Route dependent transmission costs are given by: $G_{1}=3,\;G_{2}=18$, and the sampling cost is $C_{s}=2$.
We then consider two different scenarios with no energy constraint. First of which is a scenario with three available routes where $\mathcal{N}=\{1,2,3\}$, and two different availability settings. The parameter setting for the simulations where all routes are always ON ($\boldsymbol{p}=\boldsymbol{1}$) is presented in Table I. In this table, parameters specified as intervals indicate the values that are varied along the horizontal axis of the corresponding simulation figure. For notational convenience, we define:
$\sigma_{\min}=\min_{r\in\mathcal{N}}\sigma_{r},\quad\mu_{\min}=\min_{r\in\mathcal{N}}\mu_{r},\quad\mu_{\max}=\max_{r\in\mathcal{N}}\mu_{r}.$ (73)
The parameter setting for the simulations where a single route (route $1$) is always ON is presented in Table  II where $p_{2}=p_{3}=p$ is the varying parameter.
Figure 8: Visualization of simulated optimal policies.
VIII-D Discussions
@@
Fig. 4 shows that the long-term average AoI decreases as the energy constraint is relaxed. Moreover, once $E_{\max}$ exceeds the energy consumption of the unconstrained system, the AoI converges to a constant value, as expected.
We have considered two routes that achieve approximately the same optimal AoI in the energy-unconstrained regime. A striking outcome is that, despite being six times more expensive ($G_{2}=6G_{1}$), route 2 becomes the exclusive choice of the optimal policy when $E_{\max}$ is small. The reason lies in their delay characteristics: route 1 has low delay variance, so its unconstrained optimal policy is close to zero-wait; whereas route 2 has high delay variance, so its unconstrained optimal policy naturally introduces large waiting times (long idle periods), which substantially reduce energy usage. Thus, when energy is scarce, route 2 becomes preferable.
@@
A further noteworthy observation is that the energy usage of the unconstrained joint optimal policy, $E=4.14$, exceeds that of the unconstrained optimal policies for the individual routes ($E_{1}=3.4$, $E_{2}=2.75$). This arises from the routing threshold between the two routes. Route 2 is not used in the low-delay region, which reduces the expected waiting time when it is used. This reduction in waiting time increases overall energy expenditure compared to the single-route scenarios.
@@
Fig. 5 demonstrates that MAD-Zero Wait can underperform against a simpler policy. When $p=0$, MAD-Zero Wait performs similar to the optimal policy since route $1$’s delay has a small variance in this setting. As $p$ increases, the MAD-Zero Wait policy uses the other routes, which are not suitable to zero wait policies, increasingly often. When $p=1$, MAD-Zero Wait policy is equivalent to the route $3$-Zero Wait policy. An analysis showing when we can expect MAD-Zero Wait to exhibit the behavior in Fig. 5 follows:
@@
Figures 5 and 6 demonstrate that the advantage of using the optimal policy over the MAD-Optimal policy dwindles in the low availability region (small $p$). It can also be seen that the separation between the policies happens for a larger $p$ when route $1$ has worse delay characteristics. As a result, it may be beneficial to apply the MAD-Optimal policy (smaller complexity) in cases where route $1$ is uncompetitive and routes $2,\dots,N$ have low joint availability (i.e, when $\prod_{2}^{N}1-p_{i}$ is large).
Fig. 7(a) highlights a surprising finding: a higher delay variance may counterintuitively improve the average AoI performance. In a single-route setting, reducing the delay variance typically leads to more regular update arrivals and thus lower AoI. However, this intuitive conclusion breaks down when an additional route is available, as evidenced by Fig. 7(a). Moreover, the figure shows that when $\sigma_{\min}=\sigma_{1}$ is below a certain threshold (approximately 2.5), the optimal policy actively utilizes both routes. Once this threshold is crossed, route 1 is no longer selected, and the route 2 (MAD-Optimal) with minimum delay is always used.
Fig. 7(b) demonstrates that route 1 (MDV-Optimal) provides better AoI performance than route 2 (MAD-Optimal) when $\mu_{1}=\mu_{\max}$ is relatively small. While route 2 is used in the optimal policy due to its shorter average delay, its role is marginal. Notably, the benefit of joint routing peaks when $\mu_{\max}$ is just above 3. As $\mu_{\max}$ continues to increase and exceeds approximately 4, the route 2 (MAD-Optimal) policy becomes age-optimal.
Fig. 7(c) presents the long-term average AoI values in a three-route scenario, where $\sigma_{1}=\sigma_{\min}$ is varied from 0 to 3. The optimal policy utilizes all three routes until $\sigma_{\min}$ exceeds a threshold of approximately 1.5, beyond which route 1 is no longer selected, and the policy relies solely on routes 2 and 3.Fig. 7(d) shows the long-term average AoI values for another three-route scenario, where $\mu_{3}=\mu_{\min}$ is varied from 0.7 to 1.2. In this case, all three routes consistently appear in the optimal policy, as route 3 remains the minimum delay route. However, its contribution to the overall performance becomes negligible as $\mu_{3}=\mu_{\min}$ approaches $\mu_{2}=1.2$.
Fig. 8(a) and Fig. 8(b) show threshold structure of the optimal routing decision $R_{i}^{\star}=r^{\star}(y)$ and waiting time decision $Z_{i}^{\star}=z^{\star}(y)$ when $Y_{i}=y$. The parameter configuration is specified in Table I. These results verify Theorem 1.
Overall, the proposed joint sampling and routing policy demonstrates robust improvements in AoI performance under diverse parameter settings. In particular, our simulations show that even in a basic three-route example, average AoI can be reduced by as much as $10\%$. This finding challenges conventional intuition and reveals a critical insight: routes that appear suboptimal in isolation—due to higher mean delays or variances—can meaningfully contribute to AoI minimization under a well-designed optimized handover policy.
IX Conclusion
@@ Appendix A Proof of Lemma 1
The proof closely follows the derivation in [8812616, Lemma 2], with details reproduced below.
@@
If $\lambda_{c}^{\star}\leq\lambda$,
$\exists~\pi,~\limsup_{n\to\infty}\;\frac{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[q(Y_{i},\,Z_{i},\,Y_{i+1})]}{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]}\leq\lambda.$ (75)
Moving $\lambda$ to the left-hand side yields: $\exists~\pi$
$\limsup_{n\to\infty}\frac{\frac{1}{n}\sum_{i=0}^{n-1}\Bigl(\mathbb{E}_{\pi}\bigl[q(Y_{i},\,Z_{i},\,Y_{i+1})\bigr]-\lambda\;\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]\Bigr)}{\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]}$ (76)
$\leq 0.$
Since $Y_{i}$’s over the same route are independent, the inter-sampling times $T_{i}=Y_{i}+Z_{i}$ are regenerative. Since there are $N$ routes, the expected period of the most frequently used route satisfies $\mathbb{E}[n_{k+1}-n_{k}]\leq N$, where $n_{k}$ denotes the $k$-th epoch a particular route is used. Because $T_{i}$’s are regenerative and we have $0<\mathbb{E}[D_{n_{k+1}}-D_{n_{k}}]<\infty$, for all $k$, the renewal theory [ross1995stochastic] tells us that $\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}[Z_{i}+Y_{i+1}]$ exists and is positive. Thus, there exists a policy $\pi$ such that the numerator of the left-hand side (76) is less than zero, which indicates that the infimum of the numerator in (76) is less than zero, indicating that $h(\lambda,c)\leq 0$.
@@
$\limsup_{n\to\infty}\;\frac{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[q(Y_{i},\,Z_{i},\,Y_{i+1})]}{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]}=\lambda,$ (77)
@@
$\forall~\pi,~\limsup_{n\to\infty}\;\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[q(Y_{i},\,Z_{i},\,Y_{i+1})]-\lambda\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]\geq 0,$ (78)

### v2

@@ VII Simulation Results
This section presents simulation results for practical scenarios to validate the analytical findings and evaluate the performance of our proposed algorithm. To provide a clear understanding of the system dynamics, we analyze the impact of non-linear age penalty functions, energy constraints, and stochastic route availability in isolation. This decoupled approach is adopted for the following reasons:
First, mixing strict energy constraints with other system parameters can obscure the fundamental benefits of dynamic routing. As we observe in the subsequent analysis, in low-energy regimes, the optimal policy converges to a static single-route strategy to conserve power. By isolating the energy constraint, we can distinctly characterize the transition from energy-limited static behavior to diversity-driven dynamic routing, without confounding the results with non-linear penalty dynamics.
Second, the analysis of stochastic route availability requires careful benchmarking. If the proposed policy is compared against a single-route baseline that is always available, the relative performance gain can become arbitrarily large, rendering the comparison trivial. By treating availability separately and proposing our own specific benchmarks for that case, we ensure fair and consistent comparisons against the standard always-available single-route policies established in prior literature for other constraints, allowing us to quantify the precise value added by the proposed algorithm under each unique practical constraint.
VII-A Comparing Benchmarks
@@
In this subsection, we refer to our jointly optimal sampling and routing policy as the “optimal policy” and evaluate its performance against the following benchmark policies:
@@
$\bullet$ Minimum Average Delay Routing with AoI-Optimal Sampling (MAD-Optimal): This policy always selects the route with the minimum average delay over the set of available routes at each instance. Given this selection, a modified version of the ReAVI with the minimization over the routing options is implemented to find the AoI-optimal waiting strategy and minimize the long-term average AoI. This policy follows the AoI-optimal waiting strategy as outlined in [31, Theorem 4], when $\boldsymbol{p}=\boldsymbol{1}$.
@@
$\bullet$ Minimum Average Delay Routing with Zero-Wait Sampling (MAD-Zero Wait): This policy always selects the route with the minimum average delay over the set of available routes at each instance. It is combined with a zero-wait strategy, where a new packet is sampled and transmitted immediately upon the delivery of the previous packet. The long term average AoI achieved by this policy can be analytically calculated.
@@
$\quad\quad\ +\left(\frac{\mu_{2}}{2}+\frac{\sigma_{2}^{2}}{2\mu_{2}}\right)\left(p-p^{2}\right)\mu_{2}+\left(\frac{\mu_{3}}{2}+\frac{\sigma_{3}^{2}}{2\mu_{3}}\right)p\mu_{3}.$
@@
$\bullet$ Single Route Routing with Zero-Wait Sampling (Route $X$-ZW): This policy always selects route $x$. It is combined with a zero-wait strategy.
$\bullet$ Single Route Routing with AoI-Optimal Sampling (Route $X$ Optimal): This policy always selects route $x$ and follows the AoI-optimal waiting strategy as outlined in [31, Theorem 4].
VII-B Satellite-Terrestrial Integrated Routes
@@
VII-B1 LEO Satellite Routes with Stochastic Delays
For $l\in\mathcal{N}_{\text{Sat}}$, the delay is modeled by a log-normal distribution, characterized by the following probability density function [24]:
@@ VII-B1 LEO Satellite Routes with Stochastic Delays
VII-B2 Terrestrial Routes with Stochastic Delays
If $l\in\mathcal{N}_{\text{Ter}}$, we leverage the gamma distribution to simulate the statistics of delay $y$, where the probability density function is given by [24]:
@@ VII-B2 Terrestrial Routes with Stochastic Delays
VII-C Parameter Settings
TABLE I: Simulation Parameters $E_{\max}=\infty$
@@ VII-C Parameter Settings
TABLE II: Impact of exponential growth rate $\alpha$ and sensitivity to $\mu_{1}$ variations where $Y_{\text{cap}}=50$. The table reports the average penalty cost for single-route optimal policies versus the optimal policy under the nominal $\mu_{1}$, as well as the optimal policy’s performance when $\mu_{1}$ varies by $\pm 10\%$.
$\alpha$ Nominal Case $10\%\mu_{1}$ $10\%\mu_{1}$
Route 1 Route 2 Optimal Route 1 Optimal Route 1 Optimal
0.1 1.844 1.827 1.5748 1.9627 1.63 1.7555 1.5314
0.15 11.88 11.89 9.34 14.576 10.313 9.785 8.2
0.2 9.705 9.734 6.908 14.141 8.0139 6.7628 5.457
0.3 10.026 10.072 5.812 20.381 7.2419 5.0575 3.5815
1 11.031 11.299 4.482 35.716 7.1141 3.4699 1.8937
TABLE III: Performance comparison under different penalty saturation values $Y_{\text{cap}}$ (with fixed $\alpha=\text{1}$). Parentheses denote the cost ratio relative to the optimal policy (Benchmark Cost/Optimal Cost).
$Y_{\text{cap}}$ Average Penalty Cost
Route 1 Route 2 Optimal
50 11.031 ($2.46\times$) 11.299 ($2.52\times$) 4.482
100 2.705 ($2.96\times$) 2.694 ($2.95\times$) 0.9127
200 3.244 ($3.54\times$) 3.224 ($3.52\times$) 0.9164
400 7.976 ($4.11\times$) 7.987 ($4.12\times$) 1.939
Figure 5: AoI vs. $E_{\max}$
To evaluate the performance of the proposed policy under non-linear aging costs, we adopt a truncated exponential penalty function defined as:
$f(\Delta)=a\left(\min\left(e^{\alpha\Delta},e^{\alpha Y_{\text{cap}}}\right)-1\right),$ (52)
where $\alpha>0$ determines the growth rate of the penalty, $Y_{\text{cap}}$ imposes a maximum saturation level on the cost, and $a$ is a variable scaling parameter used for normalization. We introduce the saturation threshold $Y_{\text{cap}}$ to reflect the practical reality that in many real-time systems, information effectively loses all utility beyond a certain age. Once the data staleness exceeds this critical level, any further delay does not alter the value of the penalty function.
We consider a two-route system designed to represent an “ultra-competitive” scenario where routing decisions are non-trivial. The parameters for the second route are fixed at $\mu_{2}=0.1$ and $\sigma_{2}=6.05$ with Gamma distribution, and the standard deviation of the delay of the first route is fixed at $\sigma_{1}=3.7$ with Log-normal distribution. For each distinct value of $\alpha$ tested, the mean delay of the first route, $\mu_{1}$, is tuned specifically to create a scenario where both routes are statistically competitive (i.e., offering similar average age penalties). We do not consider availability or energy restrictions in this scenario.
Finally, since $\mu_{1}$ is selected to create a high-gain scenario, we vary this parameter by $\pm 10\%$ to assess the stability of the result. This sensitivity analysis serves to verify that the reported gains are robust features of the proposed policy under the given penalty function, rather than fragile outcomes that vanish with minor deviations in system parameters.
@@
Secondly, we consider a scenario where there are $2$ available routes that are always available with the varied energy constraint $E_{\max}\in[1,5]$, under the linear age penalty function $f(\Delta)=\Delta$. The routes have the delay statistics: Log-normal distribution, $\mu_{1}=5,\sigma_{1}=1$; Gamma distribution, $\mu_{2}=1,\sigma_{2}=7.3$. Route dependent transmission costs are given by: $G_{1}=3,\;G_{2}=18$, and the sampling cost is $C_{s}=2$.
We then consider a scenario with no energy constraint with three available routes where $\mathcal{N}=\{1,2,3\}$. The parameter setting for the simulations is presented in Table  I where $p_{2}=p_{3}=p$ is the varying parameter.
VII-D Discussions
VII-D1 Impact of Penalty Parameters of Non-Linear Penalty Functions
Table II illustrates the system performance as the exponential growth rate $\alpha$ varies. The parameter $\alpha$ essentially dictates the system’s sensitivity to aging; a larger $\alpha$ imposes a steeper penalty on outdated information, thereby placing a higher premium on maintaining a low AoI. We observe that the relative benefit of the proposed joint policy becomes more pronounced in regimes with higher $\alpha$. As the penalty for high age explodes, intelligent routing becomes critical. Specifically, for $\alpha=1$, the optimal policy achieves a penalty reduction of nearly $60\%$ compared to the best single-route alternative. This confirms that in systems where stale information is costly, dynamic routing yields substantial gains over static strategies.
Furthermore, the sensitivity analysis with $\mu_{1}\pm 10\%$ demonstrates the robustness of these gains. The improvements persist even when the delay statistics deviate from the nominal ”ultra-competitive” configuration, confirming that the policy’s advantage is not an artifact of highly specific parameter tuning. A striking example is observed at $\alpha=1$ when Route 1 degrades ($\mu_{1}+10\%$). Although the stand-alone cost of Route 1 is more than three times that of Route 2, the optimal policy does not simply discard it. Instead, by opportunistically utilizing Route 1, the joint policy achieves a cost reduction of over $37\%$ compared to relying solely on the ”better” Route 2. This highlights the policy’s ability to extract value from diversity, even when route qualities are highly asymmetric.
Table III examines the effect of the saturation threshold $Y_{\text{cap}}$. This parameter represents the ”uselessness threshold” of the information; beyond this age, the data is considered equally stale, and the penalty remains constant. The results indicate a clear trend: increasing $Y_{\text{cap}}$ amplifies the gains achieved by the optimal policy. When the cap is low, the penalty function is flatter, dampening the impact of poor routing decisions. However, as $Y_{\text{cap}}$ increases, the system is exposed to the full cost of aging for longer durations. Notably, for $Y_{\text{cap}}=400$, the optimal policy reduces the average penalty by over $75\%$ compared to single-route baselines.
VII-D2 Impact of Energy Constraints
@@
Fig. 5 illustrates the impact of the energy budget $E_{\max}$ on system performance. As expected, the long-term average AoI decreases as the energy constraint is relaxed, eventually converging to a constant floor once $E_{\max}$ exceeds the consumption required for the unconstrained optimum.
@@
A noteworthy observation is that the energy consumption of the unconstrained joint optimal policy ($E=4.14$) exceeds that of the unconstrained policies for individual routes ($E_{1}=3.4$, $E_{2}=2.75$). This increase stems from the threshold-based routing structure: the joint policy avoids Route 2 in low-delay regions, effectively reducing the average waiting time required when the route is selected. This reduction in waiting time leads to a higher transmission frequency, thereby increasing the overall energy expenditure of the system compared to the single-route scenarios.
We consider two routes that yield approximately the same optimal AoI in the unconstrained regime. A striking outcome is that, despite incurring six times higher transmission cost ($G_{2}=6G_{1}$), Route 2 becomes the exclusive choice of the optimal policy when $E_{\max}$ is small. This behavior arises from the distinct delay characteristics of the routes. Route 1, characterized by low delay variance, relies on frequent updates (approaching zero-wait) to minimize age, leading to high average energy consumption. Conversely, Route 2’s high delay variance necessitates large waiting times to mitigate age penalties; these extended idle periods naturally reduce the average transmission frequency. Consequently, under strict energy constraints, the system favors Route 2 as it sustains operation at a lower average power draw.
It is also important to note that in this low-energy regime, the dynamic routing policy offers no discernible benefit over the best single-route strategy. This indicates that when energy is severely limited, the optimal strategy simplifies to the static selection of the single most viable route, as the system lacks the energy budget to exploit diversity gains.
VII-D3 Impact of Route Availability
@@
Fig. 6 demonstrates that MAD-Zero Wait can underperform against a simpler policy. When $p=0$, MAD-Zero Wait performs similar to the optimal policy since route $1$’s delay has a small variance in this setting. As $p$ increases, the MAD-Zero Wait policy uses the other routes, which are not suitable to zero wait policies, increasingly often. When $p=1$, MAD-Zero Wait policy is equivalent to the route $3$-Zero Wait policy. An analysis showing when we can expect MAD-Zero Wait to exhibit this behavior was given in subsection VII-A.
@@
Figures 6 and 7 demonstrate that the advantage of using the optimal policy over the MAD-Optimal policy dwindles in the low availability region (small $p$). It can also be seen that the separation between the policies happens for a larger $p$ when route $1$ has worse delay characteristics. As a result, it may be beneficial to apply the proposed MAD-Optimal policy (smaller complexity) in cases where the routes in $\mathcal{R}_{\infty}$ are uncompetitive and routes in $\mathcal{R}_{<\infty}$ have low joint availability (i.e, when $\big(\prod_{i\in\mathcal{R}_{<\infty}}1-p_{i}\big)$ is large).
VIII Conclusion
@@ Appendix A Proof of Lemma 1
If $\lambda_{c}^{\star}\leq\lambda$, by the definition of the optimal Lagrangian value, there exists a policy $\pi$ such that:
$\limsup_{n\to\infty}$ (54)
$\frac{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}\big[q(Y_{i},Z_{i},Y_{i+1})+c(E_{i+1}-E_{\max}(Z_{i}+Y_{i+1}))\big]}{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]}$
$\leq\lambda.$
Moving $\lambda$ to the left-hand side and combining terms over the common denominator yields: $\exists~\pi$ such that
$\limsup_{n\to\infty}\frac{\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[\Phi_{i}]}{\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]}\leq 0,$ (55)
where
$\Phi_{i}\triangleq$ $\;q(Y_{i},Z_{i},Y_{i+1})-\lambda(Z_{i}+Y_{i+1})$ (56)
$-c\big(E_{\max}(Z_{i}+Y_{i+1})-E_{i+1}\big).$
Since $Y_{i}$’s over the same route are independent, the inter-sampling times $T_{i}=Y_{i}+Z_{i}$ are regenerative. Since there are $N$ routes, the expected period of the most frequently used route satisfies $\mathbb{E}[n_{k+1}-n_{k}]\leq N$, where $n_{k}$ denotes the $k$-th epoch a particular route is used. Because $T_{i}$’s are regenerative and we have $0<\mathbb{E}[D_{n_{k+1}}-D_{n_{k}}]<\infty$, for all $k$, the renewal theory [25] tells us that $\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}[Z_{i}+Y_{i+1}]$ exists and is positive. Thus, there exists a policy $\pi$ such that the numerator of the left-hand side (55) is less than zero, which indicates that the infimum of the numerator in (55) is less than zero, indicating that $h(\lambda,c)\leq 0$.
@@
$\limsup_{n\to\infty}$ (57)
$\frac{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}\big[q(Y_{i},Z_{i},Y_{i+1})+c(E_{i+1}-E_{\max}(Z_{i}+Y_{i+1}))\big]}{\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[Z_{i}+Y_{i+1}]}$
$=\lambda,$
@@
$\forall~\pi,~\limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\pi}[\Phi_{i}]\geq 0,$ (58)